# Efficient Attention Mechanisms for Large Language Models: A Survey

### v2

@@ 2.3.1 Learning Objective
TTT [51] generalizes the meta-learning objective with different modeling architecture:
@@
Batch update tries to solve the difficulties of the training parallelism when $f_{S}$ works as a neural network. Usually, context memory is meta-learned with a batch size of 1, which is not feasible for general TTT models. Instead, analogous to chunk parallelism, TTT treats an entire chunk as a batch. There is no state updates occur within the batch (i.e., $S$ remains constant). After processing the batch, $S$ is updated once using the aggregated gradients or update signals from all samples in the batch. This strategy preserves parallel efficiency while accommodating the training requirements of more complex architectures.
@@
Titans [3] introduces momentum, which is commonly used in optimization, to strengthen the the capability of the memory update mechanism:
@@
Weight decay is another regularization technique in training, corresponding the forgetting mechanism within Linear Attention models. Gated DeltaNet [70] and Titans employs weight decay in its memory update, serving as a learned forget gate to limit the influence of very old or noisy data. It corresponds to the selective state retention mechanisms found in architectures like RetNet [52] and Mamba [17], where the decay mechanism is proven crucial for language modeling performance:
$S_{n}=\gamma_{n}S_{t-1}+\eta_{t}(v_{t}-S_{t-1}k_{t})k_{t}^{\top}$ (9)
@@ 2.4.1 Element-wise Linear Attention
Attention-Free Transformer [74] leverages a simple weight $\exp(K_{t^{\prime}}+w_{t,t^{\prime}})$ instead of $\exp(QK^{\top})$:
@@
Where $w_{t,t^{\prime}}$ is learned pair-wise position biases. Among the AFT variants, AFT-Simple remove $w_{t,t^{\prime}}$, achieving linearized inference patterns. Since the product of $K$ and $V$ is element-wise, the recurrent state size is $\mathbb{R}^{d}$ instead of outer-product state $\mathbb{R}^{d\times d}$.
RWKV [37] leverages decay mechanism on AFT-Simple. Specifically, RWKV improves AFT’s position biases with exponential decay $w_{t,i}=-(t-i)w$. The exponential formulation preserves the recurrence property while introducing the position biases.
Element-wise Linear Attention brings strong inference advantage. However, it suffers from the bottleneck of state size, under-performing matrix-based state size. Besides, even though element-wise memory is much fast than outer-product memory, the end-to-end advantage is still marginal since other components occupy more than 95% latency [52] with outer-product memory.
@@ 2.4.2 Multi-Pass Linear Attention
Gated Slot Attention [77] further introduces GLA into ABC framework [40]. Since $\tilde{K}_{n},\tilde{V}_{n}$ works as an implicit Lienar Attention, GSA improves the update as a gated form:
@@ 2.4.3 Bidirectional Linear Attention
Bidirectional attention plays an important role in encoder-style architectures such as BERT [13]. The key difference in the linear formulation between unidirectional and bidirectional attention lies in the inference bottleneck and computational pattern. Encoder-only models typically exhibit $O(N^{2})$ complexity. Moreover, each token in an encoder-only model has access to global information. As a result, bidirectional linear attention often maintains a constant-length global token pool to reduce complexity while preserving the use of the $\mathrm{softmax}$ function.
For instance, Linformer [62] reduces the number of keys and values to a constant length through an additional matrix projection. Luna [34] further extends the Linformer design by encoding the global token pool across model layers.
@@ 2.5 Hardware Implementation
When the decay is data-independent, $f_{\gamma}(\cdot)=\mathrm{const}\in(0,1]$. Note that GroupNorm [63] after Linear Attention is already a compulsory component [52], the explicit devisor of Kernelized Linear Attention in Equation 2 is unessential.
The parallel representation is simple and easy to understand but has two shortcomings. First, the parallel form still preserves the $O(N^{2})$ complexity, same as $\mathrm{softmax}$ Attention. Second, it’s complexity increases when representing the ICL-style Linear Attention in Section 2.3.
@@
While the recurrent representation reduces the computational complexity from $O(N^{2})$ to $O(N)$, it incurs substantial memory overhead during training. This is because $S_{n}$ involves storing outer products of $K_{n}$ and $V_{n}$, which is prohibitively expensive for long sequences. As a result, the recurrent form is typically restricted to the decoding stage.
@@
The chunkwise representation combines the advantages of linear complexity and hardware-friendly parallelism [20, 52]. As shown in Figure˜2(a), taking decay-style linear attention as an example, given a chunk size $B$, let $x_{[i]}$ denote the $i$-th chunk. Define the cumulative decay within a chunk as:
@@
$R_{i}=K_{[i]}^{\top}(V_{[i]}\odot\frac{\beta_{iB}}{\beta_{[i]}})+\beta_{iB}R_{i-1}$ (16)
@@
$O_{[i]}=(Q_{[i]}K_{[i]}^{\top}\odot D_{[i]})V_{[i]}+(Q_{[i]}R_{i-1})\odot\beta_{[i]}$ (17)
This formulation offers a unified view of recurrence and parallelism: the first term captures intra-chunk dependencies, while the second term propagates inter-chunk memory through a single matrix-vector product. Owing to its efficiency and parallelizability, the chunkwise representation is typically adopted during the training and prefill stages.
For ICL-style linear attention, hardware-friendly chunkwise representations have been developed using Householder transformations [71, 70]. However, for more sophisticated variants such as TTT and Titans, constructing an explicit chunkwise form remains challenging. Instead, these architectures typically rely on large batch sizes for memory updates, effectively simulating chunkwise computation through a fixed hyperparameter.
Kernel-level optimization is essential for achieving high performance. The widely adopted FLA[68] provides a Triton-based implementation for many common Linear Attention modules. Alternatively, custom implementations in CUDA or TileLang[58] are also provided by developer, which can be employed for further acceleration.
@@ 3.1 Fixed-pattern Sparse Attention
Several works exploit the structured pattern of token-level sparsity to build Fix-pattern sparsity mask for attention computation.
@@
Sparse Transformer [8] first applies local-window (row) attention, where $w$ is close to $\sqrt{N}$, then augments it with an extra column attention that summarizes previous locations and propagates information globally. GPT-3 [5] also adopts a sparse attention pattern similar to that used in Sparse Transformer.
StreamingLLM [64] found that a large amount of attention score is allocated to initial tokens in the input sequence, which they refer as “attention sink". They propose a simple Fixed-pattern Attention which keeps only the sink tokens and sliding window tokens. For instance, given an input sequence with length $n$, selected token subset $\mathcal{S}(t)$ for query token $q_{t}$ in StreamingLLM is formulated as
@@
where $s$ is the sink token size and $w$ is the sliding window size. For better hardware efficiency, StreamingLLM with block granularity [19] keeps sink tokens and local tokens in a block-wise manner, enabling efficient memory loading and computation.
@@
LongNet [14] introduces dilated attention as fixed sparse pattern for long context training and inference. Dilated Attention expands the attention field exponentially as the distance grows, thereby reducing the complexity of attention from $O(n^{2})$ to $O(n)$. Specifically, after dividing input along seqnence dimension into segments with length $w$, dilated sparse index are selected from each segments with an interval $r$. The selected indices of segment $i$ is:
@@
where $s_{i}$ denotes the denominator of the attention softmax for $O|_{r_{i},w_{i}}$. LogSparse [27] adopts an exponentially sparse attention scheme in which each position attends to only $logN$ tokens, which can be seen an instance of exponentially dilated attention.
@@ 3.2.1 Block-Sparse Attention for Prefill
Methods that use Block-Sparse Attention for prefilling approximate Top-K blocks that cover the majority of attention score with high recall ratio, thus reducing the computation complexity of attention from $O(n^{2})$ to $O(K)$.
@@
MInference [21] observed that there are there patterns in attention weights: Streaming (A Shape) Pattern, Vertical-Slash Pattern and Block-Sparse Pattern. It determines the optimal pattern for each attention head offline and dynamically builds sparse indices based on the assigned pattern during inference.
FlexPrefill [25] proposes a context-aware sparse attention mechanism that dynamically adjusts attention patterns and computational budgets in real-time.
XAttention [65] presents a block-sparse attention framework that utilizes antidiagonal scoring to predict the importance of attention blocks, which could efficiently identify and prune non-essential blocks, achieving high sparsity and substantial computational gains.
SpargeAttn [75] also employs block-level sparse attention for prefill, which is done through a double-stage online filtering process: the first stage rapidly predicts the attention map to skip certain matrix multiplications, and the second stage applies a softmax-aware filter to further eliminate unnecessary computations.
@@ 3.2.2 Block-Sparse Attention for Decode
Quest [55] approximate the criticality of each block by calculating an upper bound of attention weights. For block $K_{i}$ we maintain element-wise Min and Max Key $m_{i}$ and $M_{i}$ through
@@
DoubleSparsity [67] approximates critical tokens efficiently through reducing the matrix multiplication dimension of calculating $QK^{T}$ product. It first offline calculates outlier channels in $QK^{T}$, denoted as $C$. Then it selects Top-K tokens with highest approximate attention score $\hat{s}$ as the sparse subset $S$.
@@
ReSA [54] combines training-free block-sparse estimation and GQA sharing, contributing to better efficiency. Besides, ReSA proposes a rectification stage to control the KV cache accumulation error. ReSA shows advantage on long-sequence generation tasks.
@@ 3.2.3 Routing-based Block-Sparse Attention
SeerAttention [15, 16] train the gating network on pretrained LLMs through a self-distillation manner. To obtain the importance score for each block, it first conduct pooling to $Q$ and $K$ along the sequence dimension, denoted as $P_{k}$ and $P_{q}$. The downsampled $Q,K$ are then passed through a learnbale linear layer $W_{q}$ and $W_{k}$. Matrix multiplied results of projected $W_{q}P_{q}(Q)$ and $W_{k}P_{k}(K)$ go through the softmax operator as a gating process:
@@
Landmark [36] proposes using special landmark tokens to represent each block and trains the attention mechanism to directly retrieve Top-K blocks via these landmark tokens. However, it did not experiment on large-scale pretrained models.
MoBA [33] integrate trainable sparse attention into the pretraining stage. It proposes Mixture of Block Attention, which applies the Top-K mechanism from MoE as gating mechanism to decide critical blocks for each query token. The importance score of each block is computed by the inner product between query token $q$ and the mean pooling result of block $K_{i}$ along the token dimension:
@@
NSA [72] introduces a training-aware mix-granularity sparse attention mechanism consisting of three branches $C\in\{\text{cmp},\text{slc},\text{win}\}$, which correspond to compression, selection, and sliding window strategies, respectively. NSA leverages a differentiable compression branch to learn the block selection score.
@@
InfLLM-v2 [57] adopts a training-aware Top-K block sparse attention mechanism similar to MoBA. To Top-K block selection accuracy, it divides blocks into small-granularity kernels with overlap and performs aggregation on kernel importance scores within each block.
@@ 3.2.4 System-level Design Choices
Learning-aware Sparse Attention [72, 33, 57] begin to take kernel implementation and efficient execution into consideration. For efficient implementation of Block-Sparse Attention, FlashAttention [10] is used for attention computation in an efficient tilling mechanism, introducing requirements and opportunies for better utilization of hardware resources, including:
  • To avoid inconsistency in memory access, in SeerAttention [15] and MInference [21], block size $b$ is typically set to a relatively large value of at least 64.
  • To align with the minimal requirement of Grouped Matrix Multiplication instruction on GPU tensor cores, in NSA [72] and InfLLM-v2 [57], the number of K, V heads within a query group are set to at least 16.
  • To reduce memory access, NSA [72] and InfLLM-v2 [57] forces sharing of selected blocks among query groups, which is done through conducting pooling on block-level importance score within query groups.

### v3

@@ 2.3.1 Learning Objective
TTT [ttt2024] generalizes the meta-learning objective with different modeling architecture:
@@
Batch update tries to solve the difficulties of the training parallelism when $f_{S}$ works as a neural network. Usually, context memory is meta‑learned with a batch size of 1, which is not feasible for general TTT models. Instead, analogous to chunk parallelism, TTT treats an entire chunk as a batch. There is no state updates occur within the batch (i.e., $S$ remains consistent). After processing the batch, $S$ is updated once using the aggregated gradients or update signals from all samples in the batch. This strategy preserves parallel efficiency while accommodating the training requirements of more complex architectures.
@@
Titans [titans] introduces momentum, which is commonly used in optimization, to strengthen the the capability of the memory update mechanism:
@@
Weight decay is another regularization technique in training, corresponding the forgetting mechanism within Linear Attention models. Gated DeltaNet [gdn] and Titans employs weight decay in its memory update, serving as a learned forget gate to limit the influence of very old or noisy data. It corresponds to the selective state retention mechanisms found in architectures like RetNet [retnet] and Mamba [mamba], where the decay mechanism is proven crucial for language modeling performance:
$S_{n}=\gamma_{n}S_{t-1}+\beta_{t}(v_{t}-S_{t-1}k_{t})k_{t}^{\top}$ (9)
@@ 2.4.1 Element-wise Linear Attention
Attention-Free Transformer [aft] leverages a simple weight $\exp(K_{t^{\prime}}+w_{t,t^{\prime}})$ instead of $\exp(QK^{\top})$:
@@
Where $w_{t,t^{\prime}}$ is learned as pair-wise position biases. Among the AFT variants, AFT-Simple remove $w_{t,t^{\prime}}$, achieving linearized inference patterns. Since the product of $K$ and $V$ is element-wise, the recurrent state size is $\mathbb{R}^{d}$ instead of outer-product state $\mathbb{R}^{d\times d}$.
RWKV [rwkv] leverages decay mechanism on AFT-Simple. Specifically, RWKV improves AFT’s position biases with exponential decay $w_{t,i}=-(t-i)w$. The exponential formulation preserves the recurrence property while introducing the position biases.
Element-wise Linear Attention brings strong inference advantage. However, it suffers from the bottleneck of state size, under-performing matrix-based state size. Besides, even though element-wise memory is much fast than outer-product memory, the end-to-end advantage is still marginal since other components occupy more than 95% latency [retnet] with outer-product memory.
@@ 2.4.2 Multi-Pass Linear Attention
Gated Slot Attention [gatedslotattention] further introduces GLA into ABC framework [abc]. Since $\tilde{K}_{n},\tilde{V}_{n}$ works as an implicit Lienar Attention, GSA improves the update as a gated form:
@@ 2.4.3 Bidirectional Linear Attention
Bidirectional attention plays an important role in encoder-style architectures such as BERT [bert]. The key difference in the linear formulation between unidirectional and bidirectional attention lies in the inference bottleneck and computational pattern. Encoder-only models typically exhibit $O(N^{2})$ complexity. Moreover, each token in an encoder-only model has access to global information. As a result, bidirectional linear attention often maintains a constant-length global token pool to reduce complexity while preserving the use of the $\mathrm{softmax}$ function.
For instance, Linformer [linformer] reduces the number of keys and values to a constant length through an additional matrix projection. Luna [luna] further extends the Linformer design by encoding the global token pool across model layers.
@@ 2.5 Hardware Implementation
When the decay is data-independent, $f_{\gamma}(\cdot)=\mathrm{const}\in(0,1]$. Note that GroupNorm [groupnorm] after Linear Attention is already a compulsory component [retnet], the explicit devisor of Kernelized Linear Attention in Equation 2 is unessential.
The parallel representation is simple and easy to understand but has two shortcomings. First, the parallel form still preserves the $O(L^{2})$ complexity, same as $\mathrm{softmax}$ Attention. Second, it’s complexity increases when representing the ICL-style Linear Attention in Section 2.3.
@@
While the recurrent representation reduces the computational complexity from $O(L^{2})$ to $O(L)$, it incurs substantial memory overhead during training. This is because $S_{n}$ involves storing outer products of $K_{n}$ and $V_{n}$, which is prohibitively expensive for long sequences. As a result, the recurrent form is typically restricted to the decoding stage.
@@
The chunkwise representation combines the advantages of linear complexity and hardware-friendly parallelism [gau, retnet]. As shown in Figure˜2(a), taking decay-style linear attention as an example, given a chunk size $B$, let $x_{[i]}$ denote the $i$-th chunk. Define the cumulative decay within a chunk as:
@@
$R_{i}=K_{[i]}^{\cramped{{}^{\mathsf{T}}}}(V_{[i]}\odot\frac{\beta_{iB}}{\beta_{[i]}})+\beta_{iB}R_{i-1}$ (16)
@@
$O_{[i]}=(Q_{[i]}K_{[i]}^{\cramped{{}^{\mathsf{T}}}}\odot D_{[i]})V_{[i]}+(Q_{[i]}R_{i-1})\odot\beta_{[i]}$ (17)
This formulation offers a unified view of recurrence and parallelism: the first term captures intra-chunk dependencies, while the second term propagates inter-chunk memory through a single matrix-vector product. Owing to its efficiency and parallelizability, the chunkwise representation is typically adopted during the training and pre-filling stages.
For ICL-style linear attention, hardware-friendly chunkwise representations have been developed using Householder transformations [deltanet, gdn]. However, for more sophisticated variants such as TTT and Titans, constructing an explicit chunkwise form remains challenging. Instead, these architectures typically rely on large batch sizes for memory updates, effectively simulating chunkwise computation through a fixed hyperparameter.
Kernel-level optimization is essential for achieving high performance. The widely adopted FLA[fla] provides a Triton-based implementation for many common Linear Attention modules. Alternatively, custom implementations in CUDA or TileLang[tilelang] are also provided by developer, which can be employed for further acceleration.
@@ 3 Sparse Attention
Conceptually, Sparse Attention differs from Linear Attention in two key dimensions. From a computational perspective, Sparse Attention does not inherently reduce memory footprint, as it still requires storing the full KV cache. Instead, it primarily focuses on accelerating the generation phase by reducing the number of attended tokens. From a performance standpoint, while Linear Attention often involves lossy compression of the global context into a fixed-size state, Sparse Attention maintains the original resolution of selected tokens, allowing its theoretical upper bound to closely approach that of full attention.
@@ 3.1 Fixed-pattern Sparse Attention
Fixed-pattern methods represent the earliest and most direct approach within the Sparse Attention paradigm. Because these methods rely on static, pre-defined masks rather than data-dependent routing, they offer the highest degree of hardware friendliness and are the easiest to optimize at the infrastructure level. By leveraging predictable memory access patterns, fixed-pattern designs were the first to be successfully integrated into early long-context Transformers and modern system-level kernels.
@@
Sparse Transformer [sparsetransformer] first applies local-window (row) attention, where $w$ is close to $\sqrt{N}$, then augments it with an extra column attention that summarizes previous locations and propagates information globally. GPT-3 [gpt3] also adopts a sparse attention pattern similar to that used in Sparse Transformer.
StreamingLLM [streamingllm] found that a large amount of attention score is allocated to initial tokens in the input sequence, which they refer as “attention sink". They propose a simple Fixed-pattern Attention which keeps only the sink tokens and sliding window tokens. For instance, given an input sequence with length $n$, selected token subset $\mathcal{S}(t)$ for query token $q_{t}$ in StreamingLLM is formulated as
@@
where $s$ is the sink token size and $w$ is the sliding window size. For better hardware efficiency, StreamingLLM with block granularity [guo2024blocksparse] keeps sink tokens and local tokens in a block-wise manner, enabling efficient memory loading and computation.
@@
LongNet [longnet] introduces dilated attention as fixed sparse pattern for long context training and inference. Dilated Attention expands the attention field exponentially as the distance grows, thereby reducing the complexity of attention from $O(L^{2})$ to $O(L)$. Specifically, after dividing input along seqnence dimension into segments with length $w$, dilated sparse index are selected from each segments with an interval $r$. The selected indices of segment $i$ is:
@@
where $s_{i}$ denotes the denominator of the attention softmax for $O|_{r_{i},w_{i}}$. LogSparse [li2019enhancing] adopts an exponentially sparse attention scheme in which each position attends to only $logN$ tokens, which can be seen an instance of exponentially dilated attention.
@@ 3.2 Block Sparse Attention
The shift from fixed-pattern to blockwise sparse attention addresses the inherent rigidity of static masks, which fail to capture the dynamic, data-dependent nature of attention in complex tasks. While token-level dynamic selection offers flexibility, it often leads to irregular memory access that severely degrades GPU throughput. Consequently, block-level sparsity emerges as a superior trade-off, which provides the necessary adaptability to focus on task-relevant context while maintaining block-aligned memory access. This structured granularity allows models to leverage high-performance primitives, effectively balancing algorithmic expressiveness with system-level efficiency.
@@ 3.2.1 Block-Sparse Attention for Prefill
Methods that use Block-Sparse Attention for prefilling approximate Top-K blocks that cover the majority of attention score with high recall ratio, thus reducing the computation complexity of attention from $O(L^{2})$ to $O(LK)$.
@@
MInference [jiang2024minference] observed that there are there patterns in attention weights: Streaming (A Shape) Pattern, Vertical-Slash Pattern and Block-Sparse Pattern. It determines the optimal pattern for each attention head offline and dynamically builds sparse indices based on the assigned pattern during inference.
FlexPrefill [lai2025flexprefill] proposes a context-aware sparse attention mechanism that dynamically adjusts attention patterns and computational budgets in real-time.
XAttention [xu2025xattention] presents a block-sparse attention framework that utilizes antidiagonal scoring to predict the importance of attention blocks, which could efficiently identify and prune non-essential blocks, achieving high sparsity and substantial computational gains.
SpargeAttn [zhang2025spargeattn] also employs block-level sparse attention for prefill, which is done through a double-stage online filtering process: the first stage rapidly predicts the attention map to skip certain matrix multiplications, and the second stage applies a softmax-aware filter to further eliminate unnecessary computations.
@@ 3.2.2 Block-Sparse Attention for Decode
Quest [tang2024questqueryawaresparsityefficient] approximate the criticality of each block by calculating an upper bound of attention weights. For block $K_{i}$ we maintain element-wise Min and Max Key $m_{i}$ and $M_{i}$ through
@@
DoubleSparsity [yang2024posttrainingsparseattentiondouble] approximates critical tokens efficiently through reducing the matrix multiplication dimension of calculating $QK^{T}$ product. It first offline calculates outlier channels in $QK^{T}$, denoted as $C$. Then it selects Top-K tokens with highest approximate attention score $\hat{s}$ as the sparse subset $S$.
@@
ReSA [resa] combines training-free block-sparse estimation and GQA sharing, contributing to better efficiency. Besides, ReSA proposes a rectification stage to control the KV cache accumulation error. ReSA shows advantage on long-sequence generation tasks.
@@ 3.2.3 Routing-based Block-Sparse Attention
SeerAttention [seerattention, gao2025seerattentionrsparseattentionadaptation] train the gating network on pretrained LLMs through a self-distillation manner. To obtain the importance score for each block, it first conduct pooling to $Q$ and $K$ along the sequence dimension, denoted as $P_{k}$ and $P_{q}$. The downsampled $Q,K$ are then passed through a learnbale linear layer $W_{q}$ and $W_{k}$. Matrix multiplied results of projected $W_{q}P_{q}(Q)$ and $W_{k}P_{k}(K)$ go through the softmax operator as a gating process:
@@
DSA [deepseek3.2] inherits the self-distillation training pipeline from SeerAttention. However, to further alleviate the performance damage, DSA makes the sparsity computation on the token-level instead of block-level. Since the score computation of token-level sparsity is relatively high, DSA leverages smaller head dimension and lower precision to reduce the cost.
@@
Landmark [mohtashami2023landmark] proposes using special landmark tokens to represent each block and trains the attention mechanism to directly retrieve Top-K blocks via these landmark tokens. However, it did not experiment on large-scale pretrained models.
MoBA [moba] integrate trainable sparse attention into the pretraining stage. It proposes Mixture of Block Attention, which applies the Top-K mechanism from MoE as gating mechanism to decide critical blocks for each query token. The importance score of each block is computed by the inner product between query token $q$ and the mean pooling result of block $K_{i}$ along the token dimension:
@@
NSA [nsa] introduces a training-aware mix-granularity sparse attention mechanism consisting of three branches $C\in\{\text{cmp},\text{slc},\text{win}\}$, which correspond to compression, selection, and sliding window strategies, respectively. NSA leverages a differentiable compression branch to learn the block selection score.
@@
InfLLM-v2 [minicpm4] adopts a training-aware Top-K block sparse attention mechanism similar to MoBA. To Top-K block selection accuracy, it divides blocks into small-granularity kernels with overlap and performs aggregation on kernel importance scores within each block.
@@ 3.2.4 System-level Design Choices
Learning-aware Sparse Attention [nsa, moba, minicpm4] begin to take kernel implementation and efficient execution into consideration. For efficient implementation of Block-Sparse Attention, FlashAttention [flashattn2] is used for attention computation in an efficient tilling mechanism, introducing requirements and opportunies for better utilization of hardware resources, including:
  • To avoid inconsistency in memory access, in SeerAttention [seerattention] and MInference [jiang2024minference], block size $b$ is typically set to a relatively large value of at least 64.
  • To align with the minimal requirement of Grouped Matrix Multiplication instruction on GPU tensor cores, in NSA [nsa] and InfLLM-v2 [minicpm4], the number of K, V heads within a query group are set to at least 16.
  • To reduce memory access, NSA [nsa] and InfLLM-v2 [minicpm4] forces sharing of selected blocks among query groups, which is done through conducting pooling on block-level importance score within query groups.